OpenStudy (keke):

simplify; 2√(6) over 5√(15)

OpenStudy (anonymous):

\[2 \sqrt{2} \over 5 \sqrt{5}\]

OpenStudy (keke):

thats not the anwser

OpenStudy (anonymous):

\[2 \sqrt{10} \over 25\]

OpenStudy (anonymous):

?

OpenStudy (keke):

thanks

OpenStudy (anonymous):

\[6 \sqrt{\frac{2}{5}} \]

OpenStudy (anonymous):

Because you cancel out the \[\sqrt{3}\]'s then times by \[\sqrt{5} \over \sqrt{5}\]

OpenStudy (anonymous):

My posting is incorrect. Eliza27's is correct.